OpenStudy (anonymous):

The amount of energy released when 1 * 10^8 atoms of chlorine are coverted to cl^- ions according to equation cl + e ->cl^- is 57.86 * 10^-10 j calculate electron gain enthalpy in terms of kj/mol and eV?

OpenStudy (anonymous):

@ujjwal

OpenStudy (vishweshshrimali5):

See u have the energy released for \(1*10^8\) atoms So, u can calculate the energy released for 1 mol i.e. \(6.023*10^{23}\) this energy will be in \(J mol^{-1}\) convert it into \(kJ mol^{-1}\) using \(1 kJ = 10^3 J\) also \(1 eV = 1.602 ^ 10^{-19} J\) using this find out the energy in eV NOTE: During this last step use the energy released for \(1 atom\)

OpenStudy (anonymous):

thxx

OpenStudy (vishweshshrimali5):

:)